# Gait-Conditioned Reinforcement Learning with Multi-Phase Curriculum for Humanoid Locomotion

### v2

@@ I Introduction
Recent RL advances have enabled agile locomotion for legged robots [6]. However, reference-based methods like Adversarial Motion Priors (AMP)  [7] are constrained by dataset coverage and morphology gaps, while reference-free approaches often result in unnatural or inefficient behaviors, especially for running or transitions.
These limitations are amplified when integrating multiple gaits—standing, walking, running—into one controller. Prior approaches rely on multi-policy distillation [8], mixture-of-experts [9], or policy fusion [10], but such architectures introduce complexity, latency, and reward interference.
@@ II-A Reference-Based Locomotion Learning
Early successes in humanoid locomotion were predominantly achieved through reference-based approaches, where policies are trained to imitate curated motion capture (MoCap) data. A seminal example is AMP [7], which introduced an adversarial learning framework that encourages policies to generate human-like motions without explicit trajectory tracking. AMP has since been extended to quadrupedal locomotion [12], humanoid whole-body control [13], and even cross-morphology skill transfer such as enabling quadrupeds to learn bipedal walking behaviors [14], showcasing its versatility across robot forms.
Despite their effectiveness, reference-based methods inherently depend on large-scale, high-quality MoCap datasets and often struggle to generalize beyond the distribution of collected motions. Morphological mismatches between human demonstrators and robotic platforms—such as differences in limb proportions, joint limits, and mass distribution—can further degrade imitation performance. Additionally, controllers trained on curated motions may fail to adapt to unseen commands or unexpected environmental disturbances.
@@ II-B Reference-Free RL
To overcome these limitations, reference-free approaches have gained increasing attention, focusing on reward-driven policy learning without external demonstrations. In quadrupedal locomotion, notable examples include Learning Quadrupedal Locomotion over Challenging Terrain [15], where Central Pattern Generators (CPG) were incorporated into the control loop to promote stable rhythmic gaits, and Learning Robust Perceptive Locomotion for Quadrupedal Robots in the Wild [16], where Policies Modulating Trajectory Generators were introduced to regularize motion generation under challenging outdoor environments. These works demonstrated that careful reward engineering and structured priors can enable robust locomotion without relying on MoCap data.
In the context of bipedal robots, two main strategies have been explored. Some studies have designed fully reference-free frameworks by shaping periodic composition rewards to encourage stable walking cycles [17]. Others have incorporated predefined motion references, such as gait libraries or hybrid zero dynamics trajectories, into RL frameworks [18, 19], blending imitation and reward-driven optimization. While both approaches have achieved significant progress in sim-to-real transfer for bipedal locomotion, challenges remain in achieving fully natural, adaptable, and robust multi-gait control without relying on handcrafted motion references.
However, extending these ideas to humanoids remains challenging due to increased dynamic instability, richer contact modes, and more complex morphology.
@@ II-C Multi-Skill and Modular Architectures
Another important line of work explores how to enable a single agent to perform multiple skills or behaviors. Two widely used strategies are policy distillation [20] and DAgger (Dataset Aggregation) [21], both of which rely on supervised imitation. For example, Han et al. [22] employ multi-expert distillation to compress several task-specific controllers into a single policy capable of diverse locomotion across complex terrains. In contrast, Zhuang et al. [8] utilize DAgger to aggregate expert trajectories for robust policy training in dynamic parkour environments.
While these approaches achieve strong performance, they typically require substantial expert data generation and careful management of training curricula, which may limit their scalability to unseen behaviors or extended gait sets.
Alternatively, modular architectures have been proposed to manage multiple skills within a unified system. Mixtures of experts (MoE) frameworks [9] dynamically select among sub-policies based on context, achieving flexible behavior composition. Perpetual Humanoid Control [9] demonstrated real-time control across diverse human-like behaviors using an MoE-based controller. Multiplicative Compositional Policies (MCP) [10] introduced a multiplicative skill blending mechanism, enabling flexible recombination of low-level primitives. Although modular frameworks offer strong flexibility, they often introduce significant architectural complexity and require skill-specific supervision or expert pretraining, which can limit scalability.
@@ II-D Reward Routing and Value Function Decoupling
A complementary research direction tackles multi-behavior learning by mitigating reward interference through reward routing or decoupled value functions. MultiCritic Actor Learning [23] proposed maintaining a single actor while assigning separate critic heads for different tasks, significantly improving multi-task RL stability. Related ideas have been explored in locomotion settings as well. For instance, CPG-based or phase-conditioned policies [15, 16] combine rhythmic motion priors with RL to stabilize multi-gait behaviors.
Despite these advances, achieving coherent, natural, and efficient multi-gait transitions within a single lightweight policy—without explicit experts, motion references, or hierarchical planners—remains a substantial challenge.
@@ II-E Our Approach
In contrast, our work proposes a streamlined solution: we directly train a single recurrent policy to perform standing, walking, running, and smooth transitions within a unified, reference-free RL framework. To mitigate reward interference across gaits, we introduce a simple yet effective gait-conditioned reward routing mechanism, based on a compact gait ID encoding appended to the observation space. Furthermore, by incorporating human-inspired reward shaping derived from biomechanical insights, we promote biomechanically plausible, energy-efficient, and naturally coordinated motions without reliance on MoCap datasets, skill experts, or switching architectures.
Compared to prior modular, distillation-based, or reference-driven methods, our framework emphasizes minimal architectural overhead, unified end-to-end training, and strong generalization across diverse locomotion behaviors, offering a scalable pathway toward robust and naturalistic humanoid locomotion control.
@@ III-B Policy Architecture
Our policy adopts an asymmetric actor-critic architecture, where the actor and critic are implemented as recurrent neural networks composed of a Long Short-Term Memory (LSTM) [24] module followed by a Multi-Layer Perceptron (MLP). In an asymmetric setup [25], the actor network, designed for deployability, receives only proprioceptive observations, while the critic network is provided with additional privileged information during training to improve value estimation.
The use of an LSTM enables the policy to capture temporal dependencies across gait cycles and handle partial observability, while the MLP layers further process the temporal embeddings to output control actions or value estimates. This recurrent structure is particularly important for legged locomotion, where delayed contact effects, balance recovery, and multi-gait transitions demand memory beyond instantaneous observations.
Specifically, the actor observes proprioceptive inputs comprising the base angular velocity and gravity orientation projected into the robot’s local frame, commanded linear and angular velocities, joint position offsets from a nominal default pose, joint velocities, and the previous action taken by the policy. To enable explicit gait-mode awareness, the observation also includes a sine-cosine encoding of the gait phase and a one-hot gait ID vector indicating the current locomotion mode (e.g., standing, walking, running, or transition). These observations, which capture instantaneous and intended locomotion context, are fed into an LSTM followed by an MLP head, outputting a 23-dimensional joint position offset vector. The resulting joint targets are tracked by a joint-level proportional-derivative (PD) controller with fixed gains.
In contrast, the critic network receives an extended observation that augments the actor’s inputs with privileged information, which includes foot contact states, ground reaction forces, friction and restitution coefficients at contact points, binary contact indicators at the hip and knee joints, and externally applied disturbance forces along with their application points. This richer information enables the critic network—structured similarly with LSTM and MLP modules—to provide more stable and accurate value estimates, thus facilitating robust policy learning despite partial observability constraints on the actor.
Both actor and critic leverage temporal memory via LSTM structures, enabling reasoning over gait cycles, delayed contact effects, and transitions across gait phases, even in environments with noisy or limited sensory data.
An overview of our proposed gait-conditioned RL framework, highlighting the asymmetric actor-critic structure, gait-conditioned reward routing, and multi-phase curriculum learning strategy, is depicted in Fig. 2. The figure summarizes how gait mode information conditions both actor and critic observations, clarifying the data flow from observations through network processing, reward computation, and final action selection.
@@ III-C Gait-Conditioned Reward Routing
To support diverse locomotion behaviors within a unified policy, we introduce a gait-conditioned reward routing mechanism. At each timestep, the system classifies the current gait mode—such as standing, walking, running, or intermediate transitions (e.g., walk-to-stand, run-to-walk)—based on the commanded velocity and dynamic state. This classification is encoded as a one-hot gait ID and appended to the policy observation, conditioning the actor-critic network on the intended gait.
Each reward term is tagged to a specific gait mode or shared across modes. A gait mask is then applied during training to route only the relevant reward components to the learning objective. For instance, push-off dynamics and short contact duration are activated during running, while foot symmetry and stillness bonuses are used in standing. Transition gaits can inherit or interpolate between adjacent modes’ rewards, allowing smooth behavior blending.
As illustrated in Fig. 3, this mechanism enables independent optimization of each gait behavior without interference, while still maintaining a shared policy backbone. It stabilizes multi-gait learning, ensures clean credit assignment, and supports seamless gait switching within a single recurrent controller.
III-D Gait-Conditioned Reward Routing
@@
During training, a gait mask is applied to selectively activate mode-specific reward terms according to the current gait. As shown in Fig. 3, the complete reward vector includes shared components (e.g., task-level tracking and regulation rewards) as well as gait-related terms (e.g., contact patterns or push-off dynamics) specific to walking, running, or standing. The gait mask filters the reward vector such that only the relevant subset contributes to the learning signal.
@@
III-E Curriculum Learning Strategy
@@
  • Gait complexity: Starting from walking, we progressively add standing, running, and transitions.
  • Commanded velocity: The speed range expands from low-speed walking to high-speed running (up to $4.0$ m/s).
  • Coordination mechanisms: Reward routing, transition blending, and deceleration smoothing are gradually enabled.
We define three training phases :
  • Phase 1: Low-speed walking (002222 m/s), focusing on rhythm, foot clearance, and knee posture.
  • Phase 2: Standing and walk-to-stand transitions with emphasis on double support and postural stability.
  • Phase 3: Running at higher speeds with rewards for push-off, flight phase, and short ground contact.
This progressive curriculum improves training stability, reduces reward interference, and facilitates generalization across locomotion contexts.
III-F Human-Inspired Gait Design and Reward Shaping
@@
III-F1 Angular Momentum Compensation via Arm-Leg Coordination
@@ IV-A1 Gait Switching and Velocity Tracking
To evaluate the controller’s adaptability to varying locomotion demands, we command a velocity profile that ramps up and down between 00 and $3$ m/s. Fig. 4 (bottom) shows that the actual forward velocity closely follows the commanded profile across a range of gaits. The top subplot visualizes contact phases along with background gait mode labels.
The policy exhibits stable transitions between standing (green), walking (blue), running (orange), and transitional gaits: Walk-to-Stand (purple) and Run-to-Walk (red). These transitions occur smoothly without abrupt contact shifts, indicating robust internal gait modulation. In particular, the policy slows down naturally from running into walking (R2W), and settles into a balanced standing posture (W2S) without external resets or mode switches.
Figure 4: Top: Support contact classification with background gait labels including Stand, Walk, Run, Walk-to-Stand (W2S), and Run-to-Walk (R2W). Bottom: Commanded vs. actual forward base velocity. Gait transitions are smooth, dynamically consistent, and reflect internal coordination.
@@ V Conclusion and Future Work
Our method achieves robust and naturalistic multi-gait behaviors in physics-based simulation using a Unitree G1 humanoid. Initial real-world deployment demonstrates successful transfer of standing, walking, and walk-to-stand transitions. We are currently working to extend this to dynamic running behaviors, which are already stable in simulation.
Future directions include scaling to more complex scenarios such as traversal of uneven or deformable terrain, vision-guided locomotion in dynamic environments, and the incorporation of task-conditioned objectives for whole-body behaviors like climbing, object interaction, and human-robot collaboration.

### v3

@@ I Introduction
Recent RL advances have enabled agile locomotion for legged robots [6]. However, reference-based methods such as Adversarial Motion Priors (AMP) [7] require large-scale motion capture datasets and often struggle with morphology mismatch between human demonstrators and robot platforms. Moreover, their reliance on implicit imitation objectives limits interpretability, and prevents flexible reward design or task-specific modulation.
These limitations are amplified when integrating multiple gaits/skills into one controller. Prior approaches rely on multi-policy distillation [8], mixture-of-experts [9], or skill fusion [10], which typically require expert pretraining, switching mechanisms, and careful reward coordination across sub-policies. These architectures increase system complexity and training cost, and often suffer from interference between skill domains during deployment.
@@ II-A Reference-Based Locomotion Learning
Early successes in humanoid locomotion were largely achieved through reference-based approaches, where policies imitate curated motion capture (MoCap) data. A seminal example is AMP [7], which introduced an adversarial framework to produce human-like motions without explicit trajectory tracking. In multi-gait settings, AMP uses several motion clips (e.g., trotting, pacing) with distinct root velocity and angular velocity profiles, switching between them based on commanded velocity. A velocity-tracking reward aligns the policy’s motion with the reference clip’s kinematics, effectively mapping different velocities to specific MoCap segments. AMP has been extended to quadrupeds [12], humanoid whole-body control [13], and cross-morphology transfer such as enabling quadrupeds to walk bipedally [14].
Beyond MoCap imitation, other reference-based methods leverage analytically generated trajectories or pre-defined gait libraries. Residual learning [15, 16] refines a nominal reference by predicting additive action offsets, enhancing adaptability while retaining its structure. Guided learning directly tracks analytical references such as Hybrid Zero Dynamics (HZD)[17] or Central Pattern Generator (CPG) templates[18], embedding rhythmic priors into training.
While effective, these methods depend on high-quality MoCap or analytical trajectories and often fail to generalize beyond their distribution. Morphological mismatches—differences in limb proportions, joint limits, or mass distribution—can degrade performance, and controllers trained on narrow gait libraries may struggle with unseen commands or disturbances.
@@ II-B Reference-Free RL
Reference-free RL dispenses with predefined trajectories, instead optimizing handcrafted rewards to encourage energy-efficient, symmetric, and periodic gaits [19]. Early work [20] demonstrated bipedal gait learning via symmetry and curriculum shaping; later studies targeted stepping stone navigation [21] and achieved sim-to-real transfer on Cassie with periodic rewards [16], extended to blind stair climbing [22] and vision-guided footstep placement [23, 24]. Heightmap-based perception further broadened terrain generalization [25].
While enabling greater generalization and design freedom, reference-free methods require careful reward tuning, converge slowly, and struggle with complex maneuvers such as agile transitions or jumping.
@@ II-C Multi-Skill and Modular Architectures
Multi-skill RL aims to enable a single agent to execute diverse behaviors. Supervised strategies such as policy distillation [26] and DAgger [27] aggregate expert demonstrations to train a unified policy. For example, Han et al. [28] distill multiple task-specific controllers into a single terrain-adaptive locomotion policy, while Zhuang et al. [8] use DAgger to learn robust parkour behaviors from expert rollouts. These methods achieve strong performance but depend on extensive expert data and carefully managed curricula, limiting scalability to unseen behaviors or larger gait sets.
Modular approaches manage multiple skills via decomposed control structures. Mixture-of-Experts (MoE) controllers [9] dynamically select sub-policies, as in Perpetual Humanoid Control, which achieved real-time coordination of diverse human-like behaviors. Multiplicative Compositional Policies (MCP) [10] blend low-level primitives via multiplicative composition for flexible skill recombination. While modular frameworks offer high flexibility, they often incur significant architectural complexity and require skill-specific supervision or pretraining.
@@ II-D Reward Routing and Value Function Decoupling
A key challenge in multi-behavior RL is mitigating reward interference across heterogeneous skills. Approaches such as MultiCritic Actor Learning [29] address this by maintaining a single actor with task-specific critic heads, improving stability. In locomotion, CPG- or phase-conditioned policies [18, 30] exploit rhythmic motion priors to stabilize multi-gait control.
While prior work like DeepMimic [31] incorporates one-hot skill identifiers into the observation to differentiate motions, these are used solely for policy conditioning without altering the reward structure. In contrast, our framework employs gait IDs both for conditioning and for dynamically routing gait-specific rewards, enabling structured multi-gait learning within a unified policy.
Despite these advances, achieving smooth and efficient gait transitions without explicit experts, motion references, or hierarchical planners remains an open challenge.
@@ II-E Our Approach
Unlike AMP-based approaches [31, 7] that require large-scale motion capture datasets and often map commanded velocities to pre-recorded gait clips via velocity tracking rewards, our method operates entirely without motion references. Instead of using MoCap-derived velocity profiles to drive gait switching, we define each gait—standing, walking, running, and transitions—directly through human-inspired biomechanical reward shaping, enabling the policy to discover natural transitions without clip-based supervision. In contrast to multi-policy or distillation frameworks [26, 27] that train and merge multiple specialized experts, we employ a single recurrent policy with a compact gait-ID encoding, allowing all gaits to be learned end-to-end within one unified architecture. This design eliminates expert-switching logic, reduces model complexity, and mitigates reward interference across gaits via a simple yet effective gait-conditioned reward routing mechanism. Compared to prior modular, distillation-based, or reference-driven methods, our framework emphasizes minimal architectural overhead, unified end-to-end training, and strong generalization across diverse locomotion behaviors, offering a scalable pathway toward robust and naturalistic humanoid locomotion control.
@@ III-B Policy Architecture
Our policy adopts an asymmetric actor–critic architecture [32], with both networks implemented as Long Short-Term Memory (LSTM) [33] layers followed by Multi-Layer Perceptrons (MLPs). The actor, designed for deployment, receives only proprioceptive observations, while the critic is given additional privileged information during training to improve value estimation.
The LSTM captures temporal dependencies across gait cycles, aiding balance recovery, handling delayed contact effects, and enabling smooth multi-gait transitions under partial observability. Actor inputs include base angular velocity and gravity orientation in the local frame, commanded velocities, joint position offsets from a nominal pose, joint velocities, the previous action, a sine–cosine gait phase encoding, and a one-hot gait ID (e.g., standing, walking, running, transition). The network outputs a 23-dimensional joint position offset vector, tracked by a joint-level proportional–derivative (PD) controller with fixed gains.
The critic augments the actor’s inputs with privileged features such as foot contact states, ground reaction forces, contact friction/restitution, binary contact indicators at hip/knee joints, and externally applied disturbances with application points. This improves stability and accuracy in value estimation, even in noisy or partially observable settings.
An overview of the proposed gait-conditioned RL framework—highlighting the asymmetric architecture, gait-conditioned reward routing, and multi-phase curriculum—is shown in Fig. 2, illustrating how gait mode information conditions both actor and critic observations, from sensory input through network processing, reward computation, and final action selection.
@@ III-C Gait-Conditioned Reward Routing
During training, a gait mask is applied to selectively activate mode-specific reward terms according to the current gait. As shown in Fig. 3, the complete reward vector includes shared components (e.g., task-level tracking and regulation rewards) as well as gait-related terms (e.g., contact patterns or push-off dynamics) specific to walking, running, or standing. The gait mask filters the reward vector such that only the relevant subset contributes to the learning signal. For example, running activates push-off and short contact duration rewards; walking enables swing height and foot symmetry terms; standing focuses on upright posture and stillness bonuses while disabling locomotion-specific terms.
@@
III-D Curriculum Learning Strategy
@@
  • Gait complexity: Starting from walking, we progressively add standing, running, and transitional modes.
  • Commanded velocity: The target velocity range expands from low-speed walking to high-speed running (up to $4.0$ m/s).
  • Coordination mechanisms: Reward routing, gait-aware masking, and smooth transition blending are gradually enabled.
We train the policy in three phases:
  • Phase 1 — Walking only: Only Walk (ID = 1) is enabled. The gait ID is fixed to Walk at all timesteps; other modes are disabled, and the reward mask activates walking-specific terms only. This stage develops periodic locomotion with stable contact, adequate foot clearance, and extended knees.
  • Phase 2 — Standing and Walk to Stand (W2S): We introduce gait-ID switching and corresponding rewards. When the commanded speed norm satisfies $\|\mathbf{v}_{c}\|<0.1$ m/s, the system enters W2S (ID = 2). If low speed and double support persist for a manually specified $1.5$ s, it switches to Stand (ID = 0). This hysteresis prevents premature switching.
  • Phase 3 — Running and Run-to-Walk (R2W): Run (ID = 3) is enabled using the Froude criterion $Fr=v^{2}/(g\,l)$ with $Fr>0.5$. Upon speed reduction below this regime, the system enters R2W (ID = 4) until a manually specified $2.5$ s stable duration completes, ensuring smooth velocity decay and contact reshaping. Running- and transition-specific rewards are activated, and all five modes (Stand, Walk, W2S, Run, R2W) are co-trained with full reward routing.
This progressive curriculum improves training stability, reduces reward interference, and facilitates generalization across diverse locomotion contexts.
III-E Human-Inspired Gait Design and Reward Shaping
@@
III-E1 Angular Momentum Compensation via Arm-Leg Coordination
@@ IV-A1 Gait Switching and Velocity Tracking
To evaluate the controller’s adaptability to varying locomotion demands, we command a velocity profile that ramps up and down between 0 and $3$ m/s. Fig. 4 (bottom) shows that the actual forward velocity closely follows the commanded profile across a range of gaits. The top subplot visualizes contact phases along with background gait mode labels.
The policy exhibits stable transitions between standing (green), walking (blue), running (orange), and transitional gaits: W2S (purple) and R2W (red). These transitions occur smoothly without abrupt contact shifts, indicating robust internal gait modulation. In particular, the policy slows down naturally from running into walking , and settles into a balanced standing posture without external resets or mode switches.
Figure 4: Top: Support contact classification with background gait labels including Stand, Walk, Run, W2S, and R2W. Bottom: Commanded vs. actual forward base velocity. Gait transitions are smooth, dynamically consistent, and reflect internal coordination.
@@ IV-A2 Angular Momentum Coordination
IV-A3 Ablations: Curriculum and Reward Routing
We ablate two core components under identical architecture, hyper-parameters, and domain randomization: (i) No Curriculum: train all five gait modes (Standing, Walking, Running, W2S and R2W) jointly from scratch; (ii) No Routing: retain curriculum and gait ID in the observation, but disable reward routing so that all gait-specific terms (Table I) remain active simultaneously (mask removed). All settings are trained for an equal budget and evaluated on the same validation episodes, with results summarized in Table II.
TABLE II: Ablations under equal training budget. Len: mean episode length; Ret: mean return.
Setting Len. $\uparrow$ Ret. $\uparrow$
Ours (Curriculum + Routing) $972.6$ $89.3$
No Curriculum (5 gait modes) $31.26$ $2.45$
No Reward Routing (all terms active) $11.59$ $0.00$
Removing the curriculum forces the policy to learn all gait modes simultaneously from scratch, which greatly increases the optimization difficulty and prevents convergence within the same budget (episode length $31.26$ vs. $972.6$, return $2.45$ vs. $89.3$). Disabling reward routing, even with curriculum, activates mutually conflicting gait-specific rewards at every timestep (e.g., stillness from standing vs. forward velocity from walking/running), causing unstable gradient updates and rapid performance collapse (episode length $11.59$, return $0.00$).
@@ V Conclusion and Future Work
Our method achieves robust and naturalistic multi-gait behaviors in physics-based simulation using a Unitree G1 humanoid. Initial real-world deployment demonstrates successful transfer of standing, walking, and walk-to-stand transitions. However, gait classification currently relies on manually defined gait IDs based on commanded velocity and Froude number, which constrains the emergence of novel gait patterns. The approach also depends on extensive manual design and tuning of numerous gait-specific reward terms, making it labor-intensive and increasingly challenging for larger or more complex gait sets. Moreover, as the number of gaits grows, the shared policy may be prone to catastrophic forgetting or behavioral drift due to overlapping objectives.
We are currently working to extend this to dynamic running behaviors, which are already stable in simulation. Future directions include scaling to more complex scenarios such as traversal of uneven or deformable terrain, vision-guided locomotion in dynamic environments, and the incorporation of task-conditioned objectives for whole-body behaviors like climbing, object interaction, and human-robot collaboration.
Appendix A: Implementation Details
A.1 Policy Architecture and Training Parameters
Our policy follows a recurrent actor-critic architecture with a single-layer LSTM module. The detailed configuration used for training is summarized in Table III.
TABLE III: Policy Architecture and PPO Training Hyperparameters
Component Setting
Policy architecture LSTM + MLP
Actor hidden dimensions [32]
Critic hidden dimensions [32]
RNN type LSTM
RNN hidden size 64
RNN layers 1
Activation function ELU
Initial action noise std 0.8
Entropy coefficient 0.01
PPO iterations 30,000
All training was conducted using the rsl_rl framework with a time step of 0.02s and a command update frequency of 10Hz.
A.2 Reward Function Weights
We apply a gait-conditioned reward routing mechanism, where different subsets of reward terms are activated depending on the current gait mode. The primary reward weights used during training are listed in Table IV.
TABLE IV: Active reward terms and weights in the final multi-gait setting.
Category Reward Term Weight
Regulation rewards Vertical lin. vel. penalty ($v_{z}$) 2.0-2.0
Horizontal ang. vel. penalty ($\omega_{xy}$) 0.05-0.05
Orientation deviation 1.0-1.0
Base height penalty 10.0-10.0
DOF acceleration penalty $2.5\times 10^{-7}$
DOF velocity penalty $1\times 10^{-3}$
Collision penalty 10.0-10.0
Action rate penalty 0.01-0.01
DOF limit penalty 5.0-5.0
Alive bonus $0.15$
Hip pos. deviation 1.0-1.0
Min. torso ang. vel. $2.0$
Waist pitch deviation $1.0$
Waist roll deviation $1.0$
Waist yaw deviation $1.2$
Torso yaw smoothness $0.8$
Shoulder roll control $3.0$
Arm swing Arm–leg momentum balance $5.0$
Human-like arm swing energy $0.3$
Elbow phase tracking $2.5$
Arm swing symmetry $2.0$
Arm swing–leg amp. match $1.0$
Walking Feet swing height penalty 15.0-15.0
Contact $1.0$
Straight knee $0.1$
Feet drag penalty 0.5-0.5
Standing Contact (standing) $2.5$
Base motion (standing) $2.5$
Pose (soft upper) $4.0$
Feet alignment $0.5$
Uprightness $1.0$
Feet drag penalty 0.2-0.2
Feet flatness $2.5$
Stillness bonus $2.0$
Walk→Stand (W2S) Feet swing height penalty 20.0-20.0
Contact $1.0$
Smooth slowdown $0.1$
Feet drag penalty 0.1-0.1
Running Contact $1.0$
Feet drag penalty 1.0-1.0
Short ground contact $0.2$
Feet swing height penalty 25.0-25.0
Push-off reward $1.0$
Forward velocity reward $0.2$
Run→Walk (R2W) Contact $1.0$
Smooth slowdown $1.0$
Transition to walk speed $0.5$
Feet swing height penalty 25.0-25.0
Task rewards Tracking lin. vel. $1.5$
Tracking ang. vel. $1.0$